Drug manufactures of Drug X, which was released on the market 5 years ago to treat Crohn’s disease, have noticed that many individuals treated with the drug have developed drug-induced lupus. A cohort study was conducted for post-marketing surveillance. It has documented that 9 individuals have developed drug-induced lupus over the course of the past 5 years. If 2 individuals developed the disease at the 1-year mark, 3 developed the disease at the 2-year mark, 1 developed the disease at the 3-year mark, and the 3 individuals developed the disease at the 4-year mark, what is the incidence rate of drug-induced lupus?

A) 0.4000

B) 0.9000

C) 0.3913

D) 0.0391
Answer #1

Incidence rate of drug-induces lupus

= No. of individuals developing the disease / Total time experienced by population at risk

= 9 / No. of person years spent at risk of disease

= 9 / (2*4 + 3*3 + 1*2 + 3*1)

= 9 / 22

= 0.40 (Option A)